OpenStudy (anonymous):

Manish reaches the station at a particular time and his wife drives to the station to meet him. One day he reaches earlier and sets out on foot from the station to meet his wife on the way. He reaches home 24 min earlier than the usual time. The speed of car is 4 times than Manish’s speed on foot. How much time earlier than the usual time would he have reached home, if his wife, forewarned of his plan, had met him at the station?

OpenStudy (anonymous):

y time he reaches usually.(y-24-x) new time =24/60=2/5 so 2/5of usual time -usual time=3/5

OpenStudy (anonymous):

speed of the car=4*manish foot

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):